# Orbital magnetoelectric coupling of three dimensional Chern insulators

### v2

@@ Abstract
It worth noting that $\alpha_{zz}^{(l)}$ is only defined modulo $N_{\rm sub}Ce^{2}/h$, which is rooted in the indeterminacy of electric polarization that the choice of unit cell is arbitrary. Here $N_{\rm sub}$ is the number of inequivalent layers within the smallest unit cell. Nevertheless, the difference between $\alpha_{zz}^{(l)}$ of two neighboring layers is physically measurable and is a topological quantity. The ambiguity in the branch choice of $\alpha_{zz}^{(l)}$ is either fixed by the symmetry constraint on the total bulk response $\sum_{l}\alpha^{(l)}$ or by boundary condition at the surface. We also note that when the total number of layers increases to a certain extent, even a very small electric field applied in the $z$ direction will cause Zener tunnelling and the system then is no longer an insulator. This imposes a truncation to the otherwise diverging layer-resolved OME response.
@@
In summary, we propose a never-mentioned, quantized layer-resolved OME response in generic 3D Chern insulators. This OME response is quantized or half-quantized to $e^{2}/h$ in the presence of inversion or mirror symmetry. Notably, even in the absence of symmetries, the gradient of the OME coupling is still exactly quantized in unit of $e^{2}/h$ given by Eq. (5). We demonstrate the robustness of the quantization rule against disorder through numerical calculations and topological argument using the Streda formula. Such considerable and quantized difference between layer-resolved OME $\alpha_{zz}^{(l)}$ is in principle measurable in experiments. For example, in a slab system, applying an out-of-plane electric field would induce a quantized change in orbital magnetization at the top and bottom layers, which may be probed by SQUID-on-tip microscope [40, 41].
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (grant No. 12174257) and the National Key R & D program of China (grant No. 2020YFA0309601).
@@
I I. Expressions of layer-resolved orbital magnetization and layer Chern number
@@
I.1 Naive definition of layer-resolved orbital magnetization
@@
$\textbf{M}=\frac{\textbf{m}}{V}=\frac{1}{2V}\int{\rm d}\textbf{r}\,\textbf{r}\times\textbf{j}(\textbf{r}),$ (1)
@@
$M_{z}$ $=\frac{-e}{2A}\sum_{E_{n}<\mu}\langle\phi_{n}|\textbf{r}\times\textbf{v}|\phi_{n}\rangle\bigg{|}_{z}=-\frac{ie}{2\hbar A}\sum_{E_{n}<\mu}\langle\phi_{n}|\textbf{r}\times[H,\textbf{r}]|\phi_{n}\rangle\bigg{|}_{z}$ (2)
@@
where $H$ is the system’s Hamiltonian, $A$ is the total area, $e>0$ is the elementary charge, and $|\phi_{n}\rangle$s are the eigenstates. Here, only spinless case is discussed. The first equality comes from $\textbf{v}=\frac{i}{\hbar}[H,\textbf{r}]$. The last equality use the occupied-state projector $P$. In the following, we also need the empty-state $Q$, the complement of $P$. So, we give their definition together
$P=\sum_{E_{n}<\mu}|\phi_{n}\rangle\langle\phi_{n}|,\quad Q=1-P.$ (3)
@@
Note that this expression only applies to systems with open boundary condition (OBC). This is because position operators are ill-defined in the thermodynamic limit and terms like $\langle\phi_{n}|\textbf{r}|\phi_{n}\rangle$ diverge. This formula is particularly useful when we introduce disorder in the system in OBC. It contains both the contribution of circulation of the current density from the bulk and the contribution from the surface. Then, one can define layer-resolved orbital magnetization as
where the atomic lattice basis $|sl,\textbf{R}\rangle$ represents the site at layer $l$, sublattice $s$ of unit-cell R.
I.2 Definition of layer-resolved orbital magnetization in the thermodynamic limit
In the thermodynamic limit namely with PBC, we can derive another expression for $M_{z}$ [29] starting from Eq. (2). Since
$H=(P+Q)H(P+Q)=PHP+QHQ,$ (5)
@@ I.2 Definition of layer-resolved orbital magnetization in the thermodynamic limit
$M_{z}$ $=\frac{e}{\hbar A}{\rm Im\,Tr}\{Px(PHP+QHQ)yP\}=\frac{e}{\hbar A}{\rm Im\,Tr}\{PxPHPyP+PxQHQyP\}$ (6)
@@
${\rm Im\,Tr}\{(1-Q)xPHPy(1-Q)\}$ $={\rm Im\,Tr}\{xPHPy-QxPHPy-xPHPyQ+QxPHPyQ\}$ (7)
@@
$({\rm Tr}\{xPHPy\})^{*}={\rm Tr}\{yPHPx\}={\rm Tr}\{xyPHP\}={\rm Tr}\{yxPHP\}={\rm Tr}\{xPHPy\}\,.$ (8)
@@
$M_{z}=\frac{e}{\hbar A}{\rm Im\,Tr}\{PxQHQyP-QxPHPyQ\}.$ (9)
@@
$M_{\gamma}=\frac{e}{2\hbar A}{\rm Im}[\epsilon_{\gamma\alpha\beta}\,{\rm Tr}\{Pr_{\alpha}QHQr_{\beta}P-Qr_{\alpha}PHPr_{\beta}Q\}]$ (10)
@@
The key observation is that $P\textbf{r}Q$ and $Q\textbf{r}P$ are well-defined and regular even in an unbounded system within PBC. In principle, Eq. (9) only applies to a system which remains gapped both in bulk and at edges (since we assume $PHQ=QHP=0$), and therefore does not apply, as such, to Chern insulators. However, we can make the substitution of $H\rightarrow H-\mu$ for Chern insulators or metals [29]. Under this substitution, Eq. (9) is suitable to both systems with PBC and those with OBC.
We try to get the k-space expression [30] of Eq. (9) in a slab with PBC in the $x,y$ plane but finite in the $z$ direction, namely Eq. (3) in the main text. Suppose the eigenstates of the slab are $|\psi_{n\textbf{k}}\rangle$. First, we need to evaluate
$\langle\psi_{n^{\prime}\textbf{k}^{\prime}}|\textbf{r}|\psi_{n\textbf{k}}\rangle$ $=\int{\rm d}\textbf{r}\,e^{-i\textbf{k}^{\prime}\cdot\textbf{r}}u_{n^{\prime}\textbf{k}^{\prime}}^{*}(\textbf{r})\,\textbf{r}\,e^{i\textbf{k}\cdot\textbf{r}}u_{n\textbf{k}}(\textbf{r})$ (11)
@@
where the third equality comes from $\partial_{\textbf{k}}\langle\psi_{n^{\prime}\textbf{k}}|\psi_{n\textbf{k}}\rangle=0$. Here we choose to normalize the eigenstates over the unit cell $\Omega_{0}$, so $\langle\psi_{n\textbf{k}}|\psi_{n\textbf{k}}\rangle=N_{{\rm cell}}$, while the definition of projectors is normalized to unity ($P^{2}=P$). Therefore, we can express the projectors as
$Q\textbf{r}P$ $=\frac{1}{N_{\rm cell}}\sum_{\textbf{k},\,n\in{\rm occ.}}\frac{1}{N_{\rm cell}}\sum_{\textbf{k}^{\prime},\,n^{\prime}\in{\rm emp.}}|\psi_{n^{\prime}\textbf{k}^{\prime}}\rangle\langle\psi_{n^{\prime}\textbf{k}^{\prime}}|\textbf{r}|\psi_{n\textbf{k}}\rangle\langle\psi_{n\textbf{k}}|$ (12)
@@
Similarly,
$P\textbf{r}Q=\frac{1}{N_{\rm cell}}\sum_{n\in{\rm occ.}}\sum_{n^{\prime}\in{\rm emp.}}\sum_{\textbf{k}}|\psi_{n\textbf{k}}\rangle\,i\langle u_{n\textbf{k}}|\partial_{\textbf{k}}u_{n^{\prime}\textbf{k}}\rangle_{\Omega_{0}}\langle\psi_{n^{\prime}\textbf{k}}|\,.$ (13)
@@
$Pr_{\alpha}QHQr_{\beta}P$ $=\frac{1}{N_{\rm cell}^{2}}\sum_{\textbf{k},\textbf{k}^{\prime}}\sum_{\begin{subarray}{c}n\in{\rm occ.},\\ m\in{\rm emp.}\end{subarray}}\sum_{\begin{subarray}{c}n^{\prime}\in{\rm occ.},\\ m^{\prime}\in{\rm emp.}\end{subarray}}|\psi_{n\textbf{k}}\rangle\,i\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{m\textbf{k}}\rangle_{\Omega_{0}}\langle\psi_{m\textbf{k}}|H|\psi_{m^{\prime}\textbf{k}^{\prime}}\rangle\,i\langle u_{m^{\prime}\textbf{k}^{\prime}}|\partial_{k^{\prime}_{\beta}}u_{n^{\prime}\textbf{k}^{\prime}}\rangle_{\Omega_{0}}\langle\psi_{n^{\prime}\textbf{k}^{\prime}}|$ (14)
@@
$\epsilon_{\gamma\alpha\beta}Pr_{\alpha}QHQr_{\beta}P$ $=\epsilon_{\gamma\alpha\beta}\frac{1}{N_{\rm cell}}\sum_{\textbf{k}}\sum_{n,n^{\prime}\in{\rm occ.}}\big{(}\sum_{m,m^{\prime}\in{\rm all}}-\sum_{m,m^{\prime}\in{\rm occ.}}\big{)}$ (15)
@@
$\epsilon_{\gamma\alpha\beta}{\rm Tr}\{Pr_{\alpha}QHQr_{\beta}P\}$ $=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\sum_{m\in{\rm all}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|u_{m\textbf{k}}\rangle_{\Omega_{0}}E_{m}\langle u_{m\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle_{\Omega_{0}}$ (16)
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|H_{\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle_{\Omega_{0}}.$
@@
$\epsilon_{\gamma\alpha\beta}{\rm Tr}\{Qr_{\alpha}PHPr_{\beta}Q\}$ (17)
$=\epsilon_{\gamma\alpha\beta}\frac{1}{N_{\rm cell}^{2}}{\rm Tr}\big{\{}\sum_{\textbf{k},\textbf{k}^{\prime}}\sum_{n,n^{\prime}\in{\rm occ}}\sum_{m,m^{\prime}\in{\rm emp}}|\psi_{m\textbf{k}}\rangle i\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle\psi_{n\textbf{k}}|H|\psi_{n^{\prime}\textbf{k}^{\prime}}\rangle i\langle u_{n^{\prime}\textbf{k}^{\prime}}|\partial_{k_{\beta}}u_{m^{\prime}\textbf{k}^{\prime}}\rangle_{\Omega_{0}}\langle\psi_{m^{\prime}\textbf{k}}|\big{\}}$
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\sum_{m\in{\rm emp}}\langle\partial_{k_{\alpha}}u_{m\textbf{k}}|u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle u_{n\textbf{k}}|H_{\textbf{k}}|u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle u_{n\textbf{k}}|\partial_{k_{\beta}}u_{m\textbf{k}}\rangle_{\Omega_{0}}$
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\sum_{m\in{\rm all}}E_{n\textbf{k}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m\textbf{k}}\rangle_{\Omega_{0}}\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}$
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle E_{n\textbf{k}}=-\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle E_{n\textbf{k}}.$
@@
$|\psi_{n\textbf{k}}\rangle=\sum_{s}\sum_{l=1}^{N_{\rm L}}C_{sl,n}(\textbf{k})|sl,\textbf{k}\rangle\quad{\rm with}\quad\textbf{k}=(k_{x},k_{y}),$ (18)
@@
{$|sl,\textbf{k}\rangle&=\sum_{\textbf{R}}e^{i\textbf{k}\cdot\textbf{R}}|sl,\textbf{R}\rangle,\\ |sl,\textbf{R}\rangle&=\frac{1}{N_{\rm cell}}\sum_{\textbf{k}}e^{-i\textbf{k}\cdot\textbf{R}}|sl,\textbf{k}\rangle,\end{aligned}\right.$ (19)

### v3

@@ VI Discussions
Furthermore, in genuine 3D Chern insulator, it worth noting that $\alpha_{zz}^{(l)}$ is only defined modulo $Ce^{2}/h$, which is rooted in the indeterminacy of electric polarization that the choice of unit cell is arbitrary. Nevertheless, the difference between $\alpha_{zz}^{(l)}$ of two neighboring layers is physically measurable and is a topological quantity. The ambiguity in the branch choice of $\alpha_{zz}^{(l)}$ is either fixed by the symmetry constraint on the total bulk response $\sum_{l}\alpha^{(l)}$ or by boundary condition at the surface. We also note that when the total number of layers increases to a certain extent, even a very small electric field applied in the $z$ direction will cause Zener tunnelling and the system then is no longer an insulator. This imposes a truncation to the otherwise diverging layer-resolved OME response.
@@
In summary, we propose a never-mentioned, quantized layer-resolved OME response in generic 3D Chern insulators. This OME response is quantized or half-quantized to $e^{2}/h$ in the presence of inversion or mirror symmetry. Furthermore, even lack of symmetries, the gradient of this OME coupling is still exactly quantized in unit of $e^{2}/h$ given by Eq. (8). We further prove the quantization rule using the Streda formula, combined with numerical evidence that shows the robustness of the quantization rule against substrate effects and disorder. Through extensive first principles calculations, we propose that such considerable and quantized difference between layer-resolved OME is in principle measurable in Cr-doped $\rm{(Bi/Sb)}_{2}\rm{Te}_{3}$ [44] and $\rm{MnBi}_{2}\rm{Te}_{4}$ slabs. We also propose an detailed experimental workflow of measuring such quantization rule using state-of-the-art magnetometry by reverse propagation.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (grant No. 12174257), the National Key Research and Development Program of China (grant No. 2024YFA1410400 and 2020YFA0309601). We also thank Haibiao Zhou and Weifeng Zhi for helpful discussions.
@@
I Expressions of layer-resolved orbital magnetization and layer Chern number
@@
I.1 Naive definition of layer-resolved orbital magnetization: first approach
@@
$\textbf{M}=\frac{\textbf{m}}{V}=\frac{1}{2V}\int{\rm d}\textbf{r}\,\textbf{r}\times\textbf{j}(\textbf{r}),$ (S1)
@@
$M_{z}$ $=\frac{-e}{2A}\sum_{E_{n}<\mu}\langle\phi_{n}|\textbf{r}\times\textbf{v}|\phi_{n}\rangle\bigg{|}_{z}=-\frac{ie}{2\hbar A}\sum_{E_{n}<\mu}\langle\phi_{n}|\textbf{r}\times[H,\textbf{r}]|\phi_{n}\rangle\bigg{|}_{z}$ (S2)
@@
where $H$ is the system’s Hamiltonian, $A$ is the total area, $e>0$ is the elementary charge, and $|\phi_{n}\rangle$s are the eigenstates. Here, only spinless case is discussed. The first equality comes from $\textbf{v}=\frac{i}{\hbar}[H,\textbf{r}]$. The last equality use the occupied-state projector $P$. In the following, we also need the empty-state projector $Q$, the complement of $P$. So, we give their definition together
$P=\sum_{E_{n}<\mu}|\phi_{n}\rangle\langle\phi_{n}|,\quad Q=1-P.$ (S3)
@@
Note that this expression only applies to systems with open boundary condition (OBC). This is because position operators are ill-defined in the thermodynamic limit and terms like $\langle\phi_{n}|\textbf{r}|\phi_{n}\rangle$ diverge. It contains both the contribution of circulation of the current density from the bulk and the contribution from the surface. Then, one can define layer-resolved orbital magnetization as
$M_{z}^{(l)}=\frac{e}{\hbar A}{\rm Im}\sum_{\textbf{R},s}\langle sl,\textbf{R}|PxHyP|sl,\textbf{R}\rangle,$ (S4)
where the atomic lattice basis $|sl,\textbf{R}\rangle$ represents the site at layer $l$, sublattice $s$ of unit-cell R. This formula is particularly useful when we introduce disorder in the system in OBC.
I.2 Definition of layer-resolved orbital magnetization in the thermodynamic limit: second approach
In the thermodynamic limit namely with PBC in the $x,y$-plane, we can derive another expression for $M_{z}^{(l)}$ [32] starting from Eq. (S2). Since
$H=(P+Q)H(P+Q)=PHP+QHQ,$ (S5)
@@ I.2 Definition of layer-resolved orbital magnetization in the thermodynamic limit: second approach
$M_{z}$ $=\frac{e}{\hbar A}{\rm Im\,Tr}\{Px(PHP+QHQ)yP\}=\frac{e}{\hbar A}{\rm Im\,Tr}\{PxPHPyP+PxQHQyP\}$ (S6)
@@
${\rm Im\,Tr}\{(1-Q)xPHPy(1-Q)\}$ $={\rm Im\,Tr}\{xPHPy-QxPHPy-xPHPyQ+QxPHPyQ\}$ (S7)
@@
$({\rm Tr}\{xPHPy\})^{*}={\rm Tr}\{yPHPx\}={\rm Tr}\{xyPHP\}={\rm Tr}\{yxPHP\}={\rm Tr}\{xPHPy\}\,.$ (S8)
@@
$M_{z}=\frac{e}{\hbar A}{\rm Im\,Tr}\{PxQHQyP-QxPHPyQ\}.$ (S9)
@@
$M_{\gamma}=\frac{e}{2\hbar A}{\rm Im}[\epsilon_{\gamma\alpha\beta}\,{\rm Tr}\{Pr_{\alpha}QHQr_{\beta}P-Qr_{\alpha}PHPr_{\beta}Q\}]$ (S10)
@@
The key observation is that $P\textbf{r}Q$ and $Q\textbf{r}P$ are well-defined and regular even in an unbounded system within PBC. In principle, Eq. (S9) only applies to a system which remains gapped both in bulk and at edges (since we assume $PHQ=QHP=0$), and therefore does not apply, as such, to Chern insulators. However, we can make the substitution of $H\rightarrow H-\mu$ for Chern insulators or metals [32]. Under this substitution, Eq. (S9) is suitable to both systems with PBC and those with OBC.
We derive the k-space expression [31] of Eq. (S9) in a slab with PBC in the $x,y$ plane but finite in the $z$ direction. Suppose the eigenstates of the slab are $|\psi_{n\textbf{k}}\rangle$. First, we need to evaluate
$\langle\psi_{n^{\prime}\textbf{k}^{\prime}}|\textbf{r}|\psi_{n\textbf{k}}\rangle$ $=\int{\rm d}\textbf{r}\,e^{-i\textbf{k}^{\prime}\cdot\textbf{r}}u_{n^{\prime}\textbf{k}^{\prime}}^{*}(\textbf{r})\,\textbf{r}\,e^{i\textbf{k}\cdot\textbf{r}}u_{n\textbf{k}}(\textbf{r})$ (S11)
@@
where the third equality comes from $\partial_{\textbf{k}}\langle\psi_{n^{\prime}\textbf{k}}|\psi_{n\textbf{k}}\rangle=0$ and $N_{{\rm cell}}$ is the total number of unit cells. Here we choose to normalize the eigenstates over the unit cell’s area $\Omega_{0}$, so $\langle\psi_{n\textbf{k}}|\psi_{n\textbf{k}}\rangle=N_{{\rm cell}}$, while the definition of projectors is normalized to unity ($P^{2}=P$). Therefore, we can express the projectors as
$Q\textbf{r}P$ $=\frac{1}{N_{\rm cell}}\sum_{\textbf{k},\,n\in{\rm occ.}}\frac{1}{N_{\rm cell}}\sum_{\textbf{k}^{\prime},\,n^{\prime}\in{\rm emp.}}|\psi_{n^{\prime}\textbf{k}^{\prime}}\rangle\langle\psi_{n^{\prime}\textbf{k}^{\prime}}|\textbf{r}|\psi_{n\textbf{k}}\rangle\langle\psi_{n\textbf{k}}|$ (S12)
@@
where “occ.” and “emp.” represent indices of occupied and empty bands, respectively. Similarly,
$P\textbf{r}Q=\frac{1}{N_{\rm cell}}\sum_{n\in{\rm occ.}}\sum_{n^{\prime}\in{\rm emp.}}\sum_{\textbf{k}}|\psi_{n\textbf{k}}\rangle\,i\langle u_{n\textbf{k}}|\partial_{\textbf{k}}u_{n^{\prime}\textbf{k}}\rangle_{\Omega_{0}}\langle\psi_{n^{\prime}\textbf{k}}|\,.$ (S13)
@@
$Pr_{\alpha}QHQr_{\beta}P$ $=\frac{1}{N_{\rm cell}^{2}}\sum_{\textbf{k},\textbf{k}^{\prime}}\sum_{\begin{subarray}{c}n\in{\rm occ.},\\ m\in{\rm emp.}\end{subarray}}\sum_{\begin{subarray}{c}n^{\prime}\in{\rm occ.},\\ m^{\prime}\in{\rm emp.}\end{subarray}}|\psi_{n\textbf{k}}\rangle\,i\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{m\textbf{k}}\rangle_{\Omega_{0}}\langle\psi_{m\textbf{k}}|H|\psi_{m^{\prime}\textbf{k}^{\prime}}\rangle\,i\langle u_{m^{\prime}\textbf{k}^{\prime}}|\partial_{k^{\prime}_{\beta}}u_{n^{\prime}\textbf{k}^{\prime}}\rangle_{\Omega_{0}}\langle\psi_{n^{\prime}\textbf{k}^{\prime}}|$ (S14)
@@
$\epsilon_{\gamma\alpha\beta}Pr_{\alpha}QHQr_{\beta}P$ $=\epsilon_{\gamma\alpha\beta}\frac{1}{N_{\rm cell}}\sum_{\textbf{k}}\sum_{n,n^{\prime}\in{\rm occ.}}\big{(}\sum_{m,m^{\prime}\in{\rm all}}-\sum_{m,m^{\prime}\in{\rm occ.}}\big{)}$ (S15)
@@
$\epsilon_{\gamma\alpha\beta}{\rm Tr}\{Pr_{\alpha}QHQr_{\beta}P\}$ $=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\sum_{m\in{\rm all}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|u_{m\textbf{k}}\rangle_{\Omega_{0}}E_{m}\langle u_{m\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle_{\Omega_{0}}$ (S16)
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|H_{\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle_{\Omega_{0}}.$
@@
$\epsilon_{\gamma\alpha\beta}{\rm Tr}\{Qr_{\alpha}PHPr_{\beta}Q\}$ (S17)
$=\epsilon_{\gamma\alpha\beta}\frac{1}{N_{\rm cell}^{2}}{\rm Tr}\big{\{}\sum_{\textbf{k},\textbf{k}^{\prime}}\sum_{n,n^{\prime}\in{\rm occ.}}\sum_{m,m^{\prime}\in{\rm emp.}}|\psi_{m\textbf{k}}\rangle i\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle\psi_{n\textbf{k}}|H|\psi_{n^{\prime}\textbf{k}^{\prime}}\rangle i\langle u_{n^{\prime}\textbf{k}^{\prime}}|\partial_{k_{\beta}}u_{m^{\prime}\textbf{k}^{\prime}}\rangle_{\Omega_{0}}\langle\psi_{m^{\prime}\textbf{k}}|\big{\}}$
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\sum_{m\in{\rm emp.}}\langle\partial_{k_{\alpha}}u_{m\textbf{k}}|u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle u_{n\textbf{k}}|H_{\textbf{k}}|u_{n\textbf{k}}\rangle_{\Omega_{0}}\langle u_{n\textbf{k}}|\partial_{k_{\beta}}u_{m\textbf{k}}\rangle_{\Omega_{0}}$
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\sum_{m\in{\rm all}}E_{n\textbf{k}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m\textbf{k}}\rangle_{\Omega_{0}}\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle_{\Omega_{0}}$
$=\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\langle\partial_{k_{\beta}}u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle E_{n\textbf{k}}=-\epsilon_{\gamma\alpha\beta}\sum_{\textbf{k}}\sum_{n\in{\rm occ.}}\langle\partial_{k_{\alpha}}u_{n\textbf{k}}|\partial_{k_{\beta}}u_{n\textbf{k}}\rangle E_{n\textbf{k}}.$
@@
$|\psi_{n\textbf{k}}\rangle=\sum_{s}\sum_{l=1}^{N_{\rm L}}C_{sl,n}(\textbf{k})|sl,\textbf{k}\rangle\quad{\rm with}\quad\textbf{k}=(k_{x},k_{y}),$ (S18)
@@
{$|sl,\textbf{k}\rangle&=\sum_{\textbf{R}}e^{i\textbf{k}\cdot\textbf{R}}|sl,\textbf{R}\rangle,\\ |sl,\textbf{R}\rangle&=\frac{1}{N_{\rm cell}}\sum_{\textbf{k}}e^{-i\textbf{k}\cdot\textbf{R}}|sl,\textbf{k}\rangle,\end{aligned}\right.$ (S19)